# code-for-first-row = , code-for-last-row = , code-for-first-col = , code-for-last-col =

### v1

@@ 6 Exploratory Data Analysis
In the field of data analysis, exploratory data analysis(EDA) is the preliminary investigation of data to discover patterns, spot anomalies, test hypotheses, check assumptions, relationships and insights. A major challenge in the exploratory data analysis phase is to identify noisy data which can significantly influence its outcome. Noisy data is often characterized by a high degree of variability which cannot be explained in the dataset. In partial orderings, there is the problem of data uncertainty as a result of data noise. Noise may cause changes in the overall orderings of objects. In order to address the issue of data noise, Bruggemann & Carlsen[44] construct a probability scheme which specifies a noise-model from which a distribution for noisy values explicit expressions are derived. This faciliattes a priori estimations which determines noise level effects with respect to the order relation between any pair of objects which are affected. As such, the expected level of noise perturbations can be quantified. Bruggemann & Carlsen [58] investigate the effects of data noise or uncertainties on the partial ordering of a series of objects by adopting a fuzzy approach such that partially ordered sets are obtained as a function of noise. Based on this, it becomes possible to identify the stability range in terms of noise where the original partial order is retained. Trend analysis can be a suitable method of exploratory data analysis that can be applied in the context of posets. As shown in [4], the evolution of Hasse diagrams can be analyzed over time for the synthesis of phenomenon from which a preliminary image of the temporal trend of the phenomenon can be derived. Missing data has been identified to be a problem in non linear trend analysis [270]. A poset methodology for handling missing values is enumerated in [92] using the case study of deprivation data. The pair-wise comparison of the objects of a poset can be a useful way of dealing witth missing data. The idea is that if there is no sufficient information to conmect a given object to another then there is still the possibility to have information to make comparison with other objects in the data set [17]. Formal concept analysis is particularly suited for exploratory data analysis[323, 317] because of its human-centeredness .
@@ 7 Descriptive Data Analysis
Descriptive analysis is one of the most crucial phases of statistical data analysis and data exploration that involves summarizing and describing the primary properties of a dataset. It also involves the process of using current and historical data to identify trends and relationships between variables. In addition, it is particularly useful for communicating the evolution of data and it uses trends as the basis for further analysis to drive decision-making. Poset theory has been used in a variety of contexts for data summarization and visualization. Hasse diagrams, serve the purpose of being the partial order graphical representation which conveys a considerable amount of information on the partial order structure of the data. For this, Hasse acyclic directed graphs of posets have been used for data visualization task that correspond to: causality[241], socioeconomic analysis [88], linear models and analysis of variance[242], experimental design [243], learning analytics[244], image analytics[245], medical data analysis [78], molecular structure prediction [183, 246], spatial analysis [121, 30] multidimensional data analytics [65, 46] etc. The principal aim of a visualization tool for multidimensional and partially ordered datasets is to provide a direct representation of the data structure, reducing its complexity but retaining the essential patterns in it [88]. Several formal data analytic methods have been developed that are based on the Hasse diagram. One of such methods is the Hasse Digram Technique(HDT) introduced by Brüggemann & Voigt[35] which is essentially an application of partial order theory based on a data matrix to analyze the structure of multivariate datasets, whenever a number of options can be characterized by multiple attributes (indicators). The Hasse diagram technique or use of posets can be useful in ranking problems that involve multi-criteria assessment[27] particularly when the objects from a dataset can be viewed relative to all others. There are several key motivations for the use of Hasse Diagram Technique for ranking problems as foolows:
  • it is parameter free and as such there are no assumptions needed about the statistical distributions of the attributes.
@@
  • The presentation by a Hasse diagram avoids the arbitrariness in constructing a ranking index [31].
@@
  • Hasse diagrams allows a holistic view on all the objects which are to be ranked, without introducing artificial ranking indices [28].
@@
In [32] an alternative approach using the Hasse Diagram Technique is proposed for sorting of chemicals with respect to their potential environmental hazard. In particular, in the Hasse Diagram Technique, the dominance of one chemical over another can be established whenever all attributes can simultaneously support the dominance. Additionally, [32] also proposes the average rank method which is based on the structure of the Hasse Duagram associated to the order relations encoded in the data matrix. A disadvantage of the HDT is the absence of aggregating functions, with weights as parameters, which would allow to remove conflictive descriptor values. In order to resolve that shortcoming, Brüggemann et al [158] developed METEOR(Method of evaluation by order theory) as an extended procedure which permits to solve the dilemma by obtaining a single high-ranked object, keeping the HDT transparency and allowing “researcher participation”. METEOR was subsequently applied for the evaluation of chemicals and environmental hazards in [159, 175] . Halfon et al.[107] suggest an innovative ranking procedure using a vectorial approach for partial ordering applicable to a variety of problems in environmental toxicology. The motivation to use the poset is that it can depict both not only situations of comparison amongs chemicals but also when chemicals are incomparable with respect to each other based on environmental hazard. The ranking is visualized with the help of computer programs displaying the Hasse diagrams that are valid for both large and small datasets comprising of 34 and 8 ranked chemicals respectively. In [191] the graphical technique called POSAC is developed for multivariate data analysis in a two-dimensional space such that observations and variables can be studied simultaneously. It works by mapping the rows (e.g., states) of a matrix in a way that maximizes the preservation of their partial order with similar states located in close proximity on the map. POSAC has been demonstrated to be useful in detecting the evolution of crime analytics in large geographical areas over time.
The set of chains and the antichains of a Hasse diagram represent important structural features that can be identified from their vertical and horizontal components respectively. Bruggemann[40] found that tripartite graphs are useful in the interpretation of complex data sets by clarifying the role of indicators causing incomparabilities which are associated with the horizontal components of Hasse diagrams. In the study of the evaluation of analytic performance using partial ordering in [57], it is demonstrated that the presence of incomparabilities (being an inherent feature in partial order ranking) may disclose important characteristics. It is also shown that summary statistics of the absolute z-score, absolute skewnesss and the standard deviation from calcualtions using data on analytical performnce can be visualised on a Hasse diagram to provide a holistic view of the results.
The well-established use of lattice theory is in formal concept analysis (FCA) which provides a conceptual framework for structuring, analyzing and visualizing data, in order to make them more understandable [324]. In particular, the concept lattice with its diagram graph allows the visualization and summarization of data in a more concise representation. The main point of FCA which is of interest here is the “symmetric” view on objects and properties, whereas a Hasse diagram, as shown in only shows the order relations focusing on the objects [40].
A limitation of a Hasse diagram as a representation of a poset is as a result of its complexity when there exist a large number of edges and vertices. In this case, they are not easy to interpret as the number of vertices increases and, as noticed and they do not also provide any metric information when this is available in the original data. Cluster analysis, on the other hand, reduces the complexity of the data, but it is not designed to preserve information on comparabilities and incomparabilities. Bruggemann & Carlsen [247] combine Hasse diagrams and cluster analysis in a complexity reduction process, producing a visual output that allows the end users to jointly grasp the partial order and the metric structure of the data. This visualization process involved the following steps [88]:
@@
    Reducing dataset complexity through a clustering process based on a SOM(Self-organizing map).
@@
    Building a classical Hasse diagram on the population of clusters, that is on SOM codebooks.
@@
    Visually adding information on statistical units and clusters (particularly, information pertaining to the value of the covariates)
The use of cluster analysis and principal component analysis with HDT may be helpful in obtaining a statistically relevant data representation and in avoiding insignificant numerical differences of the attributes, which in turn would lead to insignificant comparabilities and incomparabilities and thus to very complex Hasse diagrams.[31]]
Large data sets may be difficult to deal with computationally in FCA. More especially, it is the number of formal concepts derived from a data set that is the key factor in determining if a concept lattice will be useful as a visualisation. Andrews& Orphanides [6] suggests that understandable results can be obtainable from existing data sources where this would not normally have been the case whereby formal contexts that would normally be intractable for visualisation have been processed by firstly, focusing on information of interest and secondly, reducing ‘noise’ in the context, thus revealing readable lattices that lucidly portray conceptual meaning in large data sets.
@@ 8 Applications
In this section, we summarize the data sets, evaluation methods, software and algorithmic solutions for the analysis of partially orderd data. We detail practical applications of posets in various domains of machine learning and data analysis. Since as partially ordered data and structures are ubiquitous, posets have a wide variety of applications. Although the applications are very broad, a significant number of its applications pertains to data-driven decisions in problems of sustainable development[5, 109, 4]. In this case, the main examples pertain to the analysis of complex and multidimensional systems of ordinal data and to problems of multi-criteria decision making which are relevant in social and environmental sciences. A comprehensive review of the theory behind the applications of posets in socio-econics and environmental sciences is found in [94]. The application of partial order methodology can sometimes be used as an interim process before other tools will be applied is emphasized[63]. Posets and other methods arising from lattice theory are generally useful in comparative evaluation processes [27]. It is more appropriate to use poset methodology for performing comparative evaluation than using statistical tools since as decisions on binary choices based on comparisons by applying some criteria are naturally based on order theory. Furthermore, ordering based on posets can be visualised and additional results can be derived from the mathematical theory of relations between objects. Order relations also facilitate the identification and evaluation of most relevant objects in a study as well as deriving the sequence of the importance of a criteria to perform a ranking process for ordinal data. The Hasse diagram technique remains a powerful tool to perform comparative evaluations [30]. There are several techniques for quickly identifying the minimal and maximal elements of a poset even in complex scenarios which are of fundamental interest whenever priority setting procedures are to be performed. Partially ordered sets often comprise a big amount of information on the degree of dominance among their elements and proper tools can be employed to extract and turn it into synthetic scores and rankings.
@@ 8.1 Environmental Chemistry and Toxicology
The concept of partially ordered sets and their visualisation by Hasse diagrams turns out to be very useful in many applications of environmental concern where evaluative considerations are important [107]. Environmental data have been extensively studied by Hasse Diagram Technique in [29, 30, 31, 33] with the goal to identify major relationships both among sediment samples (objects) and degradation indices (attributes). Further investigations using formal concept analysis[7] with the aim to highlight interaction among hygienic compounds and a synergism between the two toxicity tests applied to the sediment samples Lake Ontario. Hasse diagram technique[30] supports the visualization of more than two-dimensional problems to identify pollution patterns by characterizing geographical regions by their pollution with chemicals.and to suggest priority regions for further examination. The pollution pattern responsible for the location of the regions within a Hasse diagram, helps to look for remediation strategies. In alternative approach, the pollution pattern of 59 regions of Baden Wuerttemberg Germany using a methodology that combined hierarchical clustering analysis(HCA) and Hasse Diagram Technique [240] . The HCA was used for classification of the objects in the set in order to find similarity classes resulting to reduced set of object repreentatives thereby making the Hasse diagrams for analysis depicting the network structure more easily understandable. Posets can be very useful when ranking environmental chemicals by multicriteria analysis [36, 277, 9]. Posets in this case provides a solid formal framework for the ranking of objects without assigning a common scale and/or weights to the criteria. In the prediction of toxicity levels of chlorobenzene in environmental chemical polychlorinated hydrocarbons, a scheme[252] is developed based on poset theory which embeds an overall reaction network
@@ 8.2 Socio-economic Analysis
Partially ordered data are prevalent in many branches of the social and behavioral sciences. This has been driven by the fact that partial order theory employed in the applied sciences overcome the intrinsic disadvantage hidden in aggregation, if a multiple attribute system is available[10]. A simple example is response categories Agree, Neutral, Disagree, and Don’t Know, of which the first three can be ordered and the last forms a category of its own. This type of data can be derived from a wide range of applications covering different topics: multidimensional poverty, economic development, inequality measurement and many more. Partially ordered set theory has been shown to be particularly suitable to address multi-criteria decision problems since it shows where multidimensional indicator data values are expressing a conflict which can then be easily identified and resolved by appropriate corrective action [65]. The use of posets with quantitative data reduces the set of operations and choices to be made in order to synthesize indicators (normalization, aggregation), even if they are the natural representation of multidimensional ordinal data [92, 93]. Using this approach, it is possible not only to investigate the nature of the phenomenon in more detail, but also to help policy-makers in their assessments. Infact, it can be considered an efective policy tool, aimed at promoting the identifcation of disadvantaged and under-developed profiles and contexts [5]. The application of the poset-based method provides an understanding of the complexity of the evolution of a social phenomena in terms of temporal trends and comparisons between regional structures in a countrywide basis using a multideimensional data source for the process [4, 49].Incomparability captures the existence of intrinsically different “shapes” of it, so providing a more realistic picture of the phenomenon under investigation. In this respect, the impossibility to compare the profiles of different cities, due to the existence of dimensions where they perform in conflicting ways, reveals the irreducible complexity of sustainability. In census data on deprivation within regions of a country[12, 85, 86] posets has been applied since they can account for incomparabilities, which are at the basis of deprivation complexity. Partial order theory has been extensively applied to the construction of synthetic indicators in various socio-economic contexts [4, 5, 8, 3, 18, 77, 39, 59, 60, 74, 91, 93, 95, 97, 254, 255]. In the area of crime analytics[48] POSAC has been used to study structure of criminal networks in terms of the longevity, the biographical data the network sizze and so forththe The main finding is that POSAC examination of measures derived from SNA reveals[There are other citations on CRIME and Posets INclude]

### v2

@@ 6 Exploratory Data Analysis
In the field of data analysis, exploratory data analysis(EDA) is the preliminary investigation of data to discover patterns, spot anomalies, test hypothesis, check assumptions, relationships and insights. A major challenge in the exploratory data analysis phase is to identify noisy data which can significantly influence its outcome. Noisy data is often characterized by a high degree of variability which cannot be explained in the dataset. In partial orderings, there is the problem of data uncertainty[265] as a result of data noise. Noise may cause changes in the overall orderings of objects. In order to address the issue of data noise, Bruggemann & Carlsen[64] construct a probability scheme which specifies a noise-model from which explicit expressions outlining the distribution for noisy values are derived. This facilitates a priori estimations which determines noise level effects with respect to the order relation between any pair of objects which are affected. As such, the expected level of noise perturbations can be quantified. Furthermore, Carlsen & Bruggemann[79] restudy the effects of data noise or uncertainties on the partial ordering of a series of objects by adopting a fuzzy approach such that partially ordered sets are obtained as a function of noise. Based on this, it becomes possible to identify the stability range in terms of noise where the original partial order is retained. Trend analysis can be a suitable method of exploratory data analysis that can be applied in the context of posets. As shown in Alaimo et al.[1], the evolution of Hasse diagrams can be analyzed over time for the synthesis of phenomenon from which a preliminary depiction of the temporal trend of the phenomenon can be derived. Missing data has been identified to be a problem in non linear trend analysis[259]. A poset methodology for handling missing values is presented in Alaimo[125] using the case study of deprivation data. The pair-wise comparison of the objects of a poset are shown to be a useful way of dealing with missing data. The idea is that if there is no sufficient information to connect a given object to another then there is still the possibility to have information to make comparisons with other objects in the dataset[22]. In the context of formal concept analysis, it is particularly suitable for exploratory data analysis[32, 164] because of its human-centeredness.
@@ 7 Descriptive Data Analysis
Descriptive analysis is one of the most crucial phases of statistical data analysis and data exploration that involves summarizing and describing the primary properties of a dataset. It also involves the process of using current and historical data to identify trends and relationships between variables. In addition, it is particularly useful for communicating the evolution of data and it uses trends as the basis for further analysis to drive decision-making. Poset theory has been used in a variety of contexts for data summarization and visualization. Hasse diagrams serve the purpose of being the partial order graphical representation which conveys a considerable amount of information on the partial order structure of the data. For this, Hasse acyclic directed graphs of posets have been used for data visualization tasks that correspond to: causality[282], socioeconomic analysis [121], linear models and analysis of variance[24], experimental design [183], learning analytics[189], image analytics[104], medical data analysis [111], molecular structure prediction [195, 219], spatial analysis [182, 48], multidimensional analysis [103, 86, 57] , biomonitoring[285], multivariate analysis[67], sustainability analysis[88], multi-criteria decision analysis[87], environmental data analysis[284], decision support[298, 227] etc. The principal aim of a visualization tool for multidimensional and partially ordered datasets is to provide a direct representation of the data structure, reducing its complexity but retaining the essential patterns in it [121]. Several formal data analysis methods have been developed that are based on the Hasse diagram. One of such methods is the Hasse Diagram Technique(HDT) introduced by Brüggemann & Voigt[54], which is essentially an application of partial order theory based on a data matrix to analyze the structure of multivariate datasets whenever a number of options can be characterized by multiple attributes (indicators). The sorting approach using the Hasse Diagram Technique (HDT)[51], can be summarized as follows. Consider a dataset of objects to be sorted with respect to some criteria or property. Let m(i) be an appropriate attribute by which the objects are ordered. Furthermore, let $x$ and $y$ represent any two objects with $m(i,x)$ the value of the $i^{th}$ attribute of the object $x$($i=1,..,n$).Then $x\geq y$(which can mean that $x$ is evaluated better than or equal to $y$) if and only if $m(i,x)\geq m(i,y)$ for every $i=1,...,.,n.$ The HDT or use of posets can be useful in ranking problems that involves multi-criteria assessment[45] particularly when the objects from a dataset can be viewed relative to all others. There are several key motivations for the use of HDT in ranking problems which are highlighted as follows:
@@
  • Ranking using the Hasse Diagram Technique has the advantage that it can be performed without any normative constraints.
@@
In Bruggemann et al.[51] the Hasse Diagram Technique is proposed for sorting of chemicals with respect to their potential environmental hazards. In particular, in the Hasse Diagram Technique, the dominance of one chemical over another can be established whenever all attributes can simultaneously support the dominance. A disadvantage of the HDT is the absence of aggregating functions with weights as parameters which would permit the removal of conflictive descriptor values. In order to resolve this shortcoming, Simon et al [262] developed METEOR(METhod of Evaluation by ORder) as an extension of the HDT procedure using partial order theory. In contrast to the HDT method, METEOR[263] combines transparent decision support and convenient tools for data analysis with the ability to include stakeholder’s preferences in the decision process. Furthermore, METEOR solves the problem of obtaining a single high-ranked object. Another significant advantage of using METEOR is that it avoids the time consuming and resource inefficient process of trial and error selection of different descriptor priorities, since it enables the possibility to calculate the probability of having a particular linear order by descriptor prioritization[243]. METEOR has been used in the computational evaluation of chemicals and environmental hazards as presented in [291, 55]. Halfon et al.[156] suggest an innovative ranking procedure using a vectorial approach for partial ordering applicable to a variety of problems in environmental toxicology. The motivation to use the poset is that it can depict not only situations of comparison amongst chemicals but also when chemicals are incomparable with respect to each other based on environmental hazard. The ranking is then visualized with the aid of computer programs displaying the Hasse diagrams which can be applicable to both small and large datasets. The set of chains and the antichains of a Hasse diagram represent important structural features that can be identified from their vertical and horizontal components respectively. Bruggemann[60] found that tripartite graphs are useful in the interpretation of complex datasets by clarifying the role of indicators causing incomparabilities which are associated with the horizontal components of Hasse diagrams. In the study of the evaluation of analytic performance using partial ordering as described in Carlsen et al,[78], it is demonstrated that the presence of incomparabilities which are considered an inherent feature in partial order ranking may disclose important characteristics. It is also shown that summary statistics of the absolute z-score, absolute skewnesss and the standard deviation from calculations using data on analytical performnce can be visualised on a Hasse diagram to provide a holistic view of the results. In [241, 69] the graphical technique called POSAC is developed for multivariate data analysis in a two-dimensional space such that observations and variables can be studied simultaneously. It works by mapping the rows (e.g., states) of a matrix in a way that maximizes the preservation of their partial order with similar states located in close proximity on the map. POSAC has been demonstrated to be useful in detecting the evolution of crime analytics in large geographical areas over time.
The well-established use of lattice theory is in formal concept analysis (FCA) which provides a conceptual framework for structuring, analyzing and visualizing data, in order to make them more understandable [138]. In particular, the concept lattice with its Hasse graph allows the visualization and summarization of data in a more concise representation. The main characteristic of FCA which is of interest here is its “symmetric” view on objects and properties, whereas a Hasse diagram only shows the order relations which connect the objects in a graphical structure [60]. Large datasets may be difficult to deal with computationally in FCA. More especially, it is the number of formal concepts derived from a dataset that is the key factor in determining if a concept lattice will be useful for visualisation purposes. Andrews & Orphanides[9] suggests that interpretable results can be derived from existing data sources in which it would not usually have been possible. This can be explained by the fact that formal contexts which would normally be intractable for visualisation have been processed by firstly, focusing on information of interest and secondly, reducing ‘noise’ in the context, thus revealing readable lattices that precisely represent conceptual meaning in large datasets. A limitation of a Hasse diagram as a representation of a poset, primarily results from its complexity when there exist a large number of edges and vertices. In this case, they are not easily interpretable as the number of vertices increases and they do not also provide any metric information when this is available in the original data. Cluster analysis, on the other hand, reduces the complexity of the data, but it is not designed to preserve information on comparabilities and incomparabilities. Bruggemann & Carlsen [61] combine Hasse diagrams and cluster analysis in a complexity reduction process, producing a visual output that allows the end users to jointly grasp the partial order and the metric structure of the data. This visualization process involves the following steps as outlined in Fattore et al [121]:
@@
    Reducing dataset complexity through a clustering process based on a SOM(Self-Organizing Map);
@@
    Building a classical Hasse diagram on the population of clusters which are associated to the weight vectors of SOM;
@@
    Visually adding information on statistical units and clusters. In particular, information pertaining to the value of the covariates.
Alternatively, the use of cluster analysis and principal component analysis[50] with HDT may be helpful in obtaining statistically relevant data representation and in avoiding insignificant numerical differences of the attributes. As a consequence this would lead to insignificant comparabilities and incomparabilities and thus to very complex Hasse diagrams.
@@ 8 Applications
In this section, we summarize the datasets, software and algorithmic solutions for the analysis of partially ordered data. Since as most of the applications of posets in machine learning problems has already being outlined in Section 3, here we focus mostly on practical data analysis applications of posets in various domains. Partially ordered data and structures are ubiquitous, and therefore they are associated with a wide variety of applications. Although the applications are very broad, a significant number of its applications pertains to data-driven decisions in problems of sustainable development[162, 2, 160, 1]. In this case, the main examples are applicable to the analysis of complex and multidimensional systems of ordinal data and to problems of multi-criteria decision making which are relevant in socio-economic and environmental sciences[127]. The application of partial order methodology can sometimes be used as an interim process[84] before other tools are used. Posets and other methods arising from lattice theory are generally useful in comparative evaluation processes [45]. It is more appropriate to use poset methodology for performing comparative evaluation than using statistical tools since as decisions on binary choices based on comparisons by applying some criteria are naturally based on order theory. Furthermore, ordering based on posets can be visualised and additional results can be derived from the mathematical theory of relations between objects. Order relations also facilitate the identification and evaluation of most relevant objects in a study, as well as deriving the sequence of the importance of a criteria to perform a ranking process for ordinal data. The Hasse diagram technique presents a powerful tool to perform comparative evaluations [48]. A procedure has been outlined in Newlin and Patil[227] for identifying the minimal and maximal elements of a poset using their Hasse graph structures in complex scenarios, which are of fundamental interest whenever priority setting procedures are to be performed. Partially ordered sets often comprise a big amount of information on the degree of dominance among their elements and proper tools can be employed to extract and turn it into rankings for various data analysis purposes.
@@ 8.1 Environmental Chemistry and Toxicology
The concept of partially ordered sets and their visualisation by Hasse diagrams turns out to be very useful in many applications of environmental pollution data studies where evaluative considerations and assessments are important [156, 128, 73, 76, 291, 77, 266, 287, 288, 289, 49, 43]. Environmental data have been extensively studied by Hasse Diagram Technique in [47, 48, 50, 52] with the goal to identify significant relationships among sediment samples (objects) and degradation indices (attributes). Further investigations using formal concept analysis[11] with the aim to highlight interaction among hygienic compounds and a synergism between toxicity tests applied to the sediment samples from surface water sources. Hasse diagram technique[48] supports the visualization of more than two-dimensional problems to identify pollution patterns by characterizing geographical regions using their chemical pollution levels, and then suggesting priority regions for further examination. The pollution pattern which determines the location of the regions within a Hasse diagram is useful in the search for remediation strategies. In an alternative approach outlined by Restrepo & Bruggemann[242], the pollution pattern from regional studies was examined using a methodology that combined hierarchical clustering analysis(HCA) and Hasse Diagram Technique. The HCA was used for classification of the objects in the set in order to find similarity classes resulting to reduced set of object repreentatives, thereby making the Hasse diagrams for analysis depicting the network structure more easily understandable. Posets can also be very useful when ranking environmental chemicals by multicriteria analysis [55, 106, 13]. Posets in this case provides a solid formal framework for the ranking of objects without assigning a common scale or weights to the criteria. In the prediction of toxicity levels of chlorobenzene in environmental chemical polychlorinated hydrocarbons, a scheme is developed by Ivanciuc et al.[167] based on poset theory and embedded within an overall reaction network.
@@ 8.2 Socio-Economic Analysis
Partially ordered data are prevalent in many branches of the social and behavioral sciences. This has been driven by the fact that partial order theory employed in the applied sciences overcomes the intrinsic disadvantage hidden in aggregation if a multiple attribute system is available[12]. A typical example is in response categories: “Agree”, “Neutral”, “Disagree”, and “Don’t Know”, of which the first three can be ordered and the last forms a category of its own. This type of data can be derived from a wide range of applications covering different topics such as multidimensional poverty, economic development, inequality measurement etc. Partially ordered set theory has been shown to be particularly suitable to address multi-criteria decision problems since it shows where multidimensional indicator data values are expressing a conflict which can then be identified and resolved by appropriate corrective action [86]. The use of posets with quantitative data reduces the set of operations and choices to be made in order to synthesize indicators (normalization, aggregation), even if they are the natural representation of multidimensional ordinal data [125, 126]. Using this approach, it is possible not only to investigate the nature of the phenomenon in more detail, but also to help policy-makers in their assessments. Infact, it can be considered an effective policy tool aimed at promoting the identification of disadvantaged and under-developed profiles and contexts [2]. The application of the poset-based method provides an understanding of the complexity of the evolution of a social phenomena in terms of temporal trends and comparisons between regional structures in a countrywide basis using a multidimensional data source for the process [1, 71]. Incomparability captures the existence of intrinsically different forms of it, thereby providing a more realistic picture of the phenomenon under investigation. In this respect, the impracticability to compare the profiles of different cities, due to the existence of dimensions where they perform in conflicting ways, reveals the irreducible complexity of sustainability. In census data on deprivation within regions of a country, poset theory has been applied because they can account for incomparabilities which are at the basis of deprivation complexity[16, 118, 119]. Partial order theory has been extensively applied to the construction of synthetic indicators in various socio-economic contexts [1, 2, 14, 3, 23, 110, 59, 80, 81, 107, 124, 126, 130, 244, 245]. In the study of crime analytics[69], POSAC(Partially Ordered Scalogram Analysis with Coordinates) is a graphical representation, providing dimensional reduction procedures that preserve partial order. It has been used to study the structure of criminal networks in terms of their longevity, biographical data the, network size and so forth. Levy [207] performs data analysis by employing the technique of partial order analysis of crime indicators characterizing cities. The results were visualized with a scalogram based on the method of Partial Order Structuple Analysis which was developed for non-metric data analysis. Furthermore, it is noted that the approach described in Levy[207], can be useful in a broad range of problems on the stratification of cities, individuals, as well as several varieties of social indicators for classification.